OpenStudy (anonymous):

How would I find the quotient? Simplify if possible, q/7 divided by q/26

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Maybe subtract 7 from 26, I'm not sure

OpenStudy (anonymous):

Maybe multiply the two

OpenStudy (skullpatrol):

Division is a special form of multiplication.

OpenStudy (skullpatrol):

In the same way that subtraction is a special form of addition.

OpenStudy (anonymous):

Yeah, it'd make more sense to multiply than to subtract, but is that the right thing to do? Multiply 7 and 26?

OpenStudy (skullpatrol):

Regular multiplication would look like this: q/7 multiplied by q/26 = $$\frac{q}{7}*\frac{q}{26} = \frac{q^2}{7*26}$$

OpenStudy (anonymous):

q^2 ----- ? 182

OpenStudy (skullpatrol):

Correct :)

OpenStudy (anonymous):

Thank you!

OpenStudy (skullpatrol):

What we want to do here is q/7 divided by q/26 which is defined as followed:

OpenStudy (skullpatrol):

Definition of division: To divide by a nonzero number multiply by its reciprocal.

OpenStudy (skullpatrol):

Do you know what the reciprocal of $$\frac{q}{26}$$is?

OpenStudy (anonymous):

26?

OpenStudy (skullpatrol):

Interchange the numerator and the denominator.

OpenStudy (anonymous):

...oh. 26/q then lol

OpenStudy (skullpatrol):

Correct :)

OpenStudy (skullpatrol):

Now, take that number and multiply it by the other.

OpenStudy (anonymous):

26q?

OpenStudy (skullpatrol):

$$\frac{q}{7}\div\frac{q}{26}$$ Definition of division: To divide by a nonzero number multiply by its reciprocal. $$\frac{q}{7}\times\frac{26}{q}$$

OpenStudy (anonymous):

Oh my. 26q ---- 7q ?

OpenStudy (skullpatrol):

Correct, now you can simplify by "cancelling" the q's.

OpenStudy (anonymous):

26 __ 7 :D

OpenStudy (anonymous):

So what was the point in multiplying? lol

OpenStudy (skullpatrol):

By definition division is multiplying by the reciprocal. I wanted to show you what regular multiplication looked like first, before showing what multiplication by the reciprocal looked like.